A team of researchers hypothesize that watching Barney on television might lead to epilepsy in children....

  1. A team of researchers hypothesize that watching Barney on television might lead to epilepsy in children. Children with (n=150) and without seizure disorders (n=150) are compared on the basis of the hours their parents reported they spent watching Barney over a one month period. Which of the following can be calculated based on this study design. Select all that apply and outline HOW you would calculate (do not have to do actual calculations).
  1. Absolute risk for epilepsy
  2. Hazard ratios for Barney viewers
  3. Odds ratios associated with Barney viewing
  4. NNH with Barney viewing
  1. The above study:
  1. Is unethical because no child should ever be exposed to Barney
  2. Can establish a cause and effect
  3. Is limited by bias and confounders
  4. Is generalizable to all children

Justify your answer:
Answer #1

Answer:

a) we can find the absolute risk for epilepsy based on how many hours the two group of children watched the Barney and how many episodes of seizure they developed ( in the group with seizure disorder).

b)thehazard ratio (HR) is the ratio of the hazard rates corresponding to the conditions described by two levels of an explanatory variable.

hazard ratio can be calculated by dividing the no of children developed seizure episodes after watching Barney divided by the total number of children taken into the study.

c) The odds ratio iscalculated by dividing the odds of the first group by the odds in the second group.

d) The number needed to harm (NNH) is an epidemiological measure that indicates how many persons on average need to be exposed to a risk factor over a specific period to cause harm in an average of one person who would not otherwise have been harmed.

It is defined as the inverse of the absolute risk increase, and computed as {\displaystyle 1/(I_{e}-I_{u})}{\displaystyle 1/(I_{e}-I_{u})}, where {\displaystyle I_{e}}I_e is the incidence in the treated (exposed) group, and {\displaystyle I_{u}}I_{u} is the incidence in the control (unexposed) group.[1] Intuitively, the lower the number needed to harm, the worse the risk factor, with 1 meaning that every exposed person is harmed.

NNH is similar to number needed to treat(NNT), where NNT usually refers to a therapeutic intervention and NNH to a detrimental effect or a risk factor.